Find the laurent series about the indicated singularity for the function e2z/(z-1)3 at z=1


1
Expert's answer
2021-09-17T03:45:05-0400

"f(z)=\frac {e^{2z}}{(z-1)^3}" ,


Let the given complex function be analytic in an annulus "r<|z\u22121|<R"  Then "f(z)" can be expanded into Laurent's series about "z=1"


Let "z-1 =u" , then "z = u+1," Putting in "f(z)" and expanding


"f(z)=\frac {1}{u^3}[1+2(u+1)+2(u+1)^2+...]"